What is the frequency of a photon with a wavelength of 159 nm. Report your answers to three significant digits.

The frequency is = _________ s-1

1
Expert's answer
2021-02-04T07:08:05-0500

Solution:

"c=f\lambda"


"f=\dfrac{c}{\lambda}"


"f=\dfrac{299792458 (\frac{m}{s})}{159*10^{-9}}=1885487*10^9"


"f=1.89*10^{15}"